OpenStudy (yoyo2443):

A dart is thrown and hits the square target shown. The target has a circle inscribed in it, and within that circle, is inscribed another square. If the dart is equally likely to hit any point on the target, what is the probability that the dart will land on the shaded area? Not sure how I'm supposed to go about this one...

OpenStudy (yoyo2443):

The graph:

jimthompson5910 (jim_thompson5910):

Are you able to determine the radius of the circle?

OpenStudy (yoyo2443):

I'm guessing the radius is \[\sqrt{2}\]

OpenStudy (yoyo2443):

From what I see on the image anyways.

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

what is the side length of the smaller square?

OpenStudy (yoyo2443):

Not sure...

jimthompson5910 (jim_thompson5910):

do you see how the side that has label "1" is half of the side length of the smaller square?

OpenStudy (yoyo2443):

Oooh, I see it now, yeah!

jimthompson5910 (jim_thompson5910):

shaded area = (area of circle) - (area of smaller square) are you able to compute the area of the shaded region?

OpenStudy (yoyo2443):

One moment let me write it out on paper.

OpenStudy (yoyo2443):

\[A=\pi \sqrt{2}^2\] \[A=\pi 2\] \[A=6.28\] Is what I got

jimthompson5910 (jim_thompson5910):

yes 2pi or approx 6.28

jimthompson5910 (jim_thompson5910):

that's the area of the circle

jimthompson5910 (jim_thompson5910):

what's the area of the small square?

OpenStudy (yoyo2443):

\[A=2^2\] which would make it 4.

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

shaded area = (area of circle) - (area of smaller square) shaded area = 2pi - 4 ... exact area shaded area = 6.28 - 4 shaded area = 2.28 .... approx area

jimthompson5910 (jim_thompson5910):

agreed?

OpenStudy (yoyo2443):

Yep!

jimthompson5910 (jim_thompson5910):

now calculate the area of the larger square

OpenStudy (yoyo2443):

Assuming the side length is about 4..? making it 16..?

OpenStudy (robtobey2):

\[\frac{\pi \sqrt{2}^2-2^2}{\left(2 \sqrt{2}\right)^2}=\frac{1}{8} (2 \pi -4)=\frac{\pi -2}{4}=0.285398 \]

jimthompson5910 (jim_thompson5910):

`Assuming the side length is about 4..? making it 16..? ` incorrect @yoyo2443

jimthompson5910 (jim_thompson5910):

sqrt(2) makes up half the side length of the bigger square

OpenStudy (yoyo2443):

so.. \[\sqrt{2}+\sqrt{2}=2.82\] ?

jimthompson5910 (jim_thompson5910):

or 2*sqrt(2)

jimthompson5910 (jim_thompson5910):

square 2*sqrt(2) to get what?

OpenStudy (yoyo2443):

To get the side length of the bigger square, 2.82

jimthompson5910 (jim_thompson5910):

[2*sqrt(2)]^2 = 2^2*[sqrt(2)]^2 [2*sqrt(2)]^2 = 4*2 [2*sqrt(2)]^2 = 8 so the area of the bigger square is 8

jimthompson5910 (jim_thompson5910):

probability of hitting shaded region = (area of shaded region)/(area of bigger square)

OpenStudy (yoyo2443):

Ahhh okay so \[Probability= (2.28)(8)=18.24\]

jimthompson5910 (jim_thompson5910):

you divide, not multiply

OpenStudy (yoyo2443):

Oh woops, I missed the '/ '\[Probability= \frac{ 2.28 }{ 8 }=0.285\]

jimthompson5910 (jim_thompson5910):

yes that's the approximate probability

jimthompson5910 (jim_thompson5910):

the exact probability is (2pi-4)/8 = (pi-2)/4

jimthompson5910 (jim_thompson5910):

using a calculator, (pi-2)/4 = 0.28539816339744 so 0.28539816339744 is the more accurate version of the approximate answer

OpenStudy (yoyo2443):

But usually rounded down to 0.28, yeah, I got it now. Thank you very much for the help!

jimthompson5910 (jim_thompson5910):

no problem